OpenStudy (anonymous):

Subtracting Fractions: 12-3/4=

OpenStudy (kohai):

So for this one, you WILL have to make 12 an improper fraction

OpenStudy (anonymous):

12 over 1

OpenStudy (anonymous):



OpenStudy (kohai):

No that's still a fraction. I guess you don't HAVE to, but... okay, what's 1- 3/4? Or, 100-75? etc

OpenStudy (kohai):

Oh I guess you could do that. Then make them like denominators

OpenStudy (jhannybean):

Find the LCM between 1 and 4. What is that?

OpenStudy (anonymous):

4

OpenStudy (jhannybean):

Ok, so multiply the numerator and denominator \(\dfrac{12}{1}\) by 4.

OpenStudy (jhannybean):

What do you get (as a fraction)?

OpenStudy (anonymous):

and the second number is 3/4?

OpenStudy (jhannybean):

Yeah, but what do you get for the first one?

OpenStudy (nurali):

\[\frac{ 12 }{ 1 }-\frac{ 3}{ 4 }\]

OpenStudy (nurali):

LCM

OpenStudy (jhannybean):

\[\frac{12}{1} \cdot \frac{4}{4}\]

OpenStudy (nurali):

\[\frac{ 12*4-3*1 }{ 1*4 }\]

OpenStudy (anonymous):

It is 4 so then..

OpenStudy (jhannybean):

Once you get that answer, combine the two fractions as @Nurali did, and simplify.

OpenStudy (nurali):

\[\frac{ 48-3 }{ 4 }\]

OpenStudy (nurali):

\[\frac{ 45 }{ 4 }\]

OpenStudy (jhannybean):

So \(12 \cdot 4\) = 48. You have 3 ontop the other fraction. Subtract the numerators, 48- 3

OpenStudy (jhannybean):

put it over 4.

OpenStudy (anonymous):

Ok so wait.. i got confused now..